OpenStudy (anonymous):

triangle DEF is similar to triangle PQR. side DE measures 3. side EF measures 7 and side FD measures 9. side PQ measures 5x-25 and side RP measures 10x-5. find x

OpenStudy (anonymous):

Your question is unclear, so I will solve for everything. The ratio of DE to FD is equal to the ratio of PQ to RP: 3 / 9 = 5x-25 / 10x-5 3(10x-5) = 9(5x-25) 30x - 15 = 45x - 225 30x + 210 = 45x 210 = 15x 14 = x Plug in x in the expressions for the sides: 5x-25 = 5(14) - 25 = 70 - 25 = 45 10x - 5 = 10(14) - 5 = 140 - 5 = 135 Use the following proportion to solve for side QR: EF/FD = QR/RP 7 / 9 = QR / 135 9(QR) = 135(7) QR = 15(7) QR = 105 To sum up: DE = 3 EF = 7 FD = 9 PQ = 45 QR = 105 RP = 135 x = 14